# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v2

@@ 4.2 Placement for Low Node-Affinity Cluster
The key insight is that intermediate states transfers occur exclusively between corresponding layers of prefill and decoding instances. Leveraging inter-op parallelism, we group layers into stages and divide each instance into segments, termed as instance segments, with each segment maintaining one specific inter-op stage. By colocating prefill and decoding segments of the same stage within a single node, we force the transfer of intermediate states to occur only via NVLINK. Inside a node, we set the same parallelism and resource allocation for segments of the same instance. Given the typical limitation of GPUs per node (usually 8), we can enumerate possible configurations inside one node and use the simulator to identify the configurations that yield the best goodput.
As outlined in Algorithm 2, we begin by enumerating inter-op parallelism degrees to get all the possible instance segments. For each segment, we get all possible intra-node configurations by calling get_intra_node_configs. Then we use simulation to find the optimal one and replicate it to satisfy the target traffic rate.
@@ 4.3 Online scheduling
Combat busrtiness. Burstiness in workloads can cause a deluge of KV caches to transfer from prefill to decoding instances, risking memory overload on decoding instances. To circumvent this, DistServe employs a “pull” method for KV cache transmission rather than a “push” approach – decoding instances fetch KV cache from prefill instances as needed, using the GPU memory of prefill instances as a queuing buffer. Hence, each type of instance operates at its own pace without complex coordination.
@@
DistServe does not implement advanced runtime policies like preemption [23] and fault tolerance [49], which are complementary to disaggregation. Nevertheless, we discuss how they fit into DistServe. In DistServe, the FCFS policy can lead to a “convoy effect”, where longer requests block shorter ones in the prefill stage. Incorporating preemptive strategies, as suggested in existing literature [44], could enhance efficiency and is feasible within our system’s architecture. While not a primary focus in the current DistServe, fault tolerance is a critical aspect for consideration. In traditional colocation- and replication-based systems, a fault in one instance typically does not disrupt other replica instances. However, in DistServe, the dependency between prefill and decoding instances introduces the risk of fault propagation. For example, a fault in a single decoding instance mapped to multiple prefill instances could potentially cripple the entire service and cluster. We leave both as future work.
@@ 5 Implementation
DistLLM is an end-to-end distributed serving system for LLMs with a placement algorithm module, a RESTful API frontend, an orchestration layer, and a parallel execution engine. The algorithm module, frontend, and orchestration layer are implemented with 6.5K lines of Python code. The parallel execution engine is implemented with 8.1K lines of C++/CUDA code.
The placement algorithm module implements the algorithm and the simulator mentioned in §4 which gives the placement decision for a specific model and cluster setting. The frontend supports OpenAI API compatible interface where clients can specify the sampling parameters like maximum output length and temperature. The orchestration layer manages the prefill and decoding instances, responsible for request dispatching, KV caches transmission, and results delivery. It utilizes NCCL [5] for cross-node GPU communication and asynchronous cudaMemcpy for intra-node communication, which avoids blocking the GPU during transmission. Each instance is powered by a parallel execution engine, which uses Ray [30] actor to implement GPU workers that execute the LLM inference and manage the KV Cache in a distributed manner. It integrates many recent LLM optimizations like continuous batching [45], FlashAttention [18], PagedAttention [28] and supports popular open-source LLMs such as OPT [47] and LLaMA [43].
@@ 6 Evaluation
In this section, we evaluate DistServe under different sizes of LLMs ranging from 13B to 175B and various application datasets including chatbot, code-completion, and summarization. The evaluation shows that DistServe consistently outperforms the current state-of-the-art system across all the settings (§6.2). Specifically, DistServe can handle up to 4.48×4.48$\times$4.48 × higher rates and 10.2×10.2$\times$10.2 × more stringent SLO while meeting the latency requirements for over 90% requests. Additionally, we analyze the latency breakdown in DistServe to show the communication overhead is insubstantial thanks to our bandwidth-aware placement algorithm (§6.3) and do ablation studies of our techniques (§6.4).
@@
Chatbot OPT-13B 26GB 0.2s 0.1s ShareGPT [7]
Chatbot OPT-66B 132GB 0.4s 0.1s ShareGPT [7]
@@ 6.1 Experiments Setup
Model and workloads setup. Similar to prior work on LLM serving [28], we choose the OPT [47] series, which is a representative LLM family widely used in academia and industry. We use FP16 precision in all experiments. For workloads, as shown in Table 1, We choose three typical LLM applications and set the SLOs empirically based on their service target because there exists no available SLO settings for these applications as far as we know. For each application, we select a suitable dataset and sample requests from it for evaluation. Since all the datasets do not include timestamps, we generate request arrival times using Poisson distribution with different request rates. Due to the space limit, we test the chatbot workload on all three OPT models and the other two workloads on OPT-66B, which matches the largest size in the recent open-source LLM series [43].
@@
  • Chatbot [1]: We use the ShareGPT dataset [7] for the chatbot application, which is a collection of user-shared conversations with ChatGPT. For OPT-13B, the TTFT SLO is set to 0.2s for responsiveness and the TPOT SLO is set to 0.1s which is higher than the normal human read speed. For OPT-66B and OPT-175B, we slightly relax the two SLOs due to the increase of model execution latency.
@@
  • Code completion [12]: We use the HumanEval [12] dataset for the code completion task. It includes 164 programming problems with a function signature or docstring which is commonly used in academia to evaluate code completion models. Since the code completion tool is used as a personal real-time coding assistant, we set both SLOs to be stringent.
@@
  • Summarization [4]: It is a popular LLM task to generate a concise summary for a long article, essay, or even an academic paper. We use LongBench [11] dataset which contains the summarization task. As shown in Figure 7, LongBench has much longer input lengths than the other two datasets. So we set a loose TTFT SLO but require a stringent TPOT.
@@
Metrics. We use SLO attainment as the major evaluation metric. Under a specific SLO attainment goal (say, 90%), we are concerned with two things: the maximum per-GPU goodput and the minimal SLO the system can handle. We are particularly interested in an SLO attainment of 90% (indicated by the vertical lines in all curve plots), but will also vary the rate and latency requirements to observe how the SLO attainment changes. To accurately understand the respective impacts of the two latency requirements on the system, we also present the proportion of requests that only meet one of these SLOs.
@@
Baseline. We compare DistServe to the state-of-the-art serving system vLLM [28]. It supports iteration-level scheduling proposed by Orca [45] and PagedAttention to reduce memory fragmentation caused by KV cache allocation. However, it colocates and batches the prefill and decoding computation to maximize the overall system throughput and struggles to meet the latency requirements in a cost-efficient way. Since vLLM only supports intra-op parallelism, we follow previous work [28] to set intra-op equals 1, 4, and 8 for the three OPT models, respectively.
@@ 6.2 End-to-end Experiments
In this Section, we compare the end-to-end performance of DistServe against vLLM on real application datasets.
@@
Chatbot. We evaluate the performance of DistServe on the chatbot application for all three OPT models. The first row of Figure 8 illustrates that when we gradually increase the rate, more requests will violate the latency requirements and the SLO attainment decreases. The vertical line shows the maximum per-GPU rate the system can handle to meet latency requirements for over 90% of the requests. The dotted and dashed lines show the achieved SLO attainment for only TTFT or TPOT requirements, respectively.
@@
On the ShareGPT dataset, DistServe can sustain 2.0×2.0$\times$2.0 ×3.41×3.41$\times$3.41 × higher request rate compared to vLLM. This is because DistLLM eliminates the prefill-decoding interference through disaggregation. Two phases can optimize their own objectives by allocating different resources and employing tailored parallelism strategies. As a result, the gap between the curve that only meets TTFT requirements (Dist-TTFT) and the one that only meets TPOT requirements (Dist-TPOT) is relatively small. Specifically, by analyzing the chosen placement strategy [^3] All the placements chosen by DistServe can be found in Appendix B. for 175B, we find the prefill instance has inter-op = 3, intra-op = 3; and the decoding instance has inter-op = 3, intra-op = 4. Under this placement, DistServe can effectively balance the load between the two instances on ShareGPT, meeting latency requirements at the lowest cost. This non-trivial placement strategy is challenging to manually find, proving the effectiveness of the algorithm. In the case of vLLM, collocating prefill and decoding greatly slows down the decoding phase, thereby significantly increasing TPOT. Due to the stringent TPOT requirements of chatbot applications, although vLLM meets the TTFT SLO for most requests, the overall SLO attainment is dragged down by a large number of requests that violate the TPOT SLO.
@@
The second row of Figure 8 indicates the robustness to the changing latency requirements of the two systems. We fix the rate and then linearly scale the two latency requirements in Table 1 simultaneously using a parameter called SLO Scale. As SLO Scale decreases, the latency requirement is more stringent. We aim to observe the most stringent SLO Scale that the system can withstand while still achieving the attainment target. Figure 8 shows that DistServe can achieve 1.4×1.4$\times$1.4 ×1.8×1.8$\times$1.8 × more stringent SLO than vLLM, thus providing more engaging service quality to the users.
@@
Code completion. Figure 9(a) shows the performance of DistServe on the code completion task when serving OPT-66B. DistServe can sustain 3.2×3.2$\times$3.2 × higher request rate and 1.5×1.5$\times$1.5 × more stringent SLO than vLLM. As a real-time coding assistant, the code completion task demands lower TTFT than chatbot, this leads to both systems ultimately being constrained by the TTFT requirement. However, in comparison, by eliminating the interference of the decoding jobs and automatically increasing intra-operation parallelism in prefill instances through the searching algorithm, DistServe reduces the average latency of the prefill jobs, thereby meeting the TTFT requirements of more requests.
@@
Summarization. Figure 9(b) shows the performance of DistServe on the summarization task when serving OPT-66B. DistServe achieves 4.48×4.48$\times$4.48 × higher request rate and 10.2×10.2$\times$10.2 × more stringent SLO than vLLM. The requests sampled from LongBench dataset have long input lengths, which brings significant pressure to the prefill computation. However, due to the loose requirement of TTFT for the summarization task, the TPOT service quality becomes particularly important. The vLLM, which collocates prefill and decoding phases, with long prefill jobs, experiences a greater slowdown in the decoding phase and fails to meet the TPOT requirement.
@@
Figure 10: Left: Latency breakdown when serving OPT-175B on ShareGPT dataset with DistServe. Right: The CDF function of KV Cache transmission time for OPT models.
@@ 6.3 Latency Breakdown
Figure 10(a) shows the latency breakdown for the OPT-175B models on ShareGPT dataset. We chose OPT-175B because the KV Cache transmission is more demanding for larger models. In fact, even for OPT-175B, the KV Cache transmission only accounts for less than 0.1% of the total latency. Even by examining the CDF of the absolute transmission time shown in Figure 10(b), we observe that over 95% of requests experience a delay of less than 30ms, despite our testbed having only limited cross-node bandwidth. This is due to the algorithm described in §4.2, where we require the prefill and decoding instance to maintain the same stage on one machine, enabling the use of intra-node NVLINK bandwidth for transmission, thus significantly reducing transmission delay.
@@ 6.4 Ablation Studies
Results. Figure 11 shows the performance of the four systems when serving OPT-13B on ShareGPT dataset. "vLLM++" has the same performance as "vLLM" because we find the default non-parallelism setting has the best per-GPU goodput. This further demonstrates the importance of disaggregation. The interference between the prefill and decoding phases significantly reduces the potential performance improvement through adjusting parallelism. In contrast, "DistLLM-High" can achieve further improvements over "DistLLM-Low" because it is not constrained by the deployment constraint that the prefill and decoding instance on one node should share the same model stage. Through disaggregation, we can use tailored parallelism strategies for prefill and decoding instances and optimize their targets without the coupling effects.

### v3

@@ 4.2 Placement for Low Node-Affinity Cluster
The key insight is that KV cache transfer occurs exclusively between corresponding layers of prefill and decoding instances. Leveraging inter-op parallelism, we group layers into stages and divide each instance into segments, termed as instance segments, with each segment maintaining one specific inter-op stage. By colocating prefill and decoding segments of the same stage within a single node, we force the transfer of intermediate states to occur only via NVLINK. Inside a node, we set the same parallelism and resource allocation for segments of the same instance. Given the typical limitation of GPUs per node (usually 8), we can enumerate possible configurations inside one node and use the simulator to identify the configurations that yield the best goodput.
As outlined in Algorithm 2, we begin by enumerating inter-op parallelism degrees to get all the possible instance segments. For each segment, we get all possible intra-node parallelism configurations by calling get_intra_node_configs. Then we use simulation to find the optimal one and replicate it to satisfy the target traffic rate.
@@ 4.3 Online scheduling
Combat busrtiness. Burstiness in workloads can cause a deluge of KV caches to transfer from prefill to decoding instances, risking memory overload on decoding instances. To circumvent this, DistServe employs a “pull” method for KV cache transmission rather than a “push” approach – decoding instances fetch KV cache from prefill instances as needed, using the GPU memory of prefill instances as a queuing buffer. This way, the prefill instance can continue handling other prefill jobs by simply retaining the KV Cache in the GPU memory after processing the prompt. Hence, each type of instance operates at its own pace without complex coordination.
Figure 6: DistServe Runtime System Architecture
@@
Preemption and fault tolerance. DistServe does not implement advanced runtime policies like preemption [26] and fault tolerance [58], which are complementary to disaggregation. Nevertheless, we discuss how they fit into DistServe. In DistServe, the FCFS policy can lead to a “convoy effect”, where longer requests block shorter ones in the prefill stage. Incorporating preemptive strategies, as suggested in existing literature [53], could enhance efficiency and is feasible within our system’s architecture. While not a primary focus in the current DistServe, fault tolerance is a critical aspect for consideration. In traditional colocation- and replication-based systems, a fault in one instance typically does not disrupt other replica instances. However, in DistServe, the dependency between prefill and decoding instances introduces the risk of fault propagation. For example, a fault in a single decoding instance mapped to multiple prefill instances could potentially cripple the entire service and cluster. We leave both as future work.
@@ 5 Implementation
DistServe is an end-to-end distributed serving system for LLMs with a placement algorithm module, a RESTful API frontend, an orchestration layer, and a parallel execution engine. The algorithm module, frontend, and orchestration layer are implemented with 6.5K lines of Python code. The parallel execution engine is implemented with 8.1K lines of C++/CUDA code.
The placement algorithm module implements the algorithm and the simulator mentioned in §4 which gives the placement decision for a specific model and cluster setting. The frontend supports an OpenAI API-compatible interface where clients can specify the sampling parameters like maximum output length and temperature. The orchestration layer manages the prefill and decoding instances, responsible for request dispatching, KV cache transmission, and results delivery. It utilizes NCCL [6] for cross-node GPU communication and asynchronous CudaMemcpy for intra-node communication, which avoids blocking the GPU computation during transmission. Each instance is powered by a parallel execution engine, which uses Ray [35] actor to implement GPU workers that execute the LLM inference and manage the KV Cache in a distributed manner. It integrates many recent LLM optimizations like continuous batching [54], FlashAttention [20], PagedAttention [32] and supports popular open-source LLMs such as OPT [56] and LLaMA [51].
@@ 6 Evaluation
In this section, we evaluate DistServe under different sizes of LLMs ranging from 13B to 175B and various application datasets including chatbot, code-completion, and summarization. The evaluation shows that DistServe consistently outperforms the current state-of-the-art system across all the settings (§6.2). Specifically, DistServe can handle up to 7.4×7.4$\times$7.4 × higher rates and 12.6×12.6$\times$12.6 × more stringent SLO while meeting the latency requirements for over 90% requests. Additionally, we analyze the latency breakdown in DistServe to show the communication overhead is insubstantial thanks to our bandwidth-aware placement algorithm (§6.3) and do ablation studies of our techniques (§6.4). Finally, we profile the execution time of our placement algorithm (§6.5).
@@
Chatbot OPT-13B 26GB 0.25s 0.1s ShareGPT [8]
Chatbot OPT-66B 132GB 2.5s 0.15s ShareGPT [8]
@@ 6.1 Experiments Setup
Model and workloads setup. Similar to prior work on LLM serving [32], we choose the OPT [56] model series, which is a representative LLM family widely used in academia and industry. Newer GPT model families are adopting memory-efficient attention mechanisms like GQA [10] and MQA [44]. DistServe will show better performance on these models because the transmission overhead is lower due to the decrease in KV cache size. We choose OPT which uses the classic MHA [52] to put enough pressure on the transmission overhead. We use FP16 precision in all experiments. For workloads, as shown in Table 1, We choose three typical LLM applications and set the SLOs empirically based on their service target because there exists no available SLO settings for these applications as far as we know. For each application, we select a suitable dataset and sample requests from it for evaluation. Since all the datasets do not include timestamps, we generate request arrival times using Poisson distribution with different request rates. Due to the space limit, we test the chatbot workload on all three OPT models and the other two workloads on OPT-66B, which matches the largest size in the recent open-source LLM series [51].
@@
  • Chatbot [1]: We use the ShareGPT dataset [8] for the chatbot application, which is a collection of user-shared conversations with ChatGPT. For OPT-13B, the TTFT SLO is set to 0.25s for responsiveness and the TPOT SLO is set to 0.1s which is higher than the normal human read speed. For OPT-66B and OPT-175B, we slightly relax the two SLOs due to the increase in model execution latency.
@@
  • Code completion [14]: We use the HumanEval [14] dataset for the code completion task. It includes 164 programming problems with a function signature or docstring which is used to evaluate the performance of code completion models. Since the code completion model is used as a personal real-time coding assistant, we set both SLOs to be stringent.
@@
  • Summarization [5]: It is a popular LLM task to generate a concise summary for a long article, essay, or even an academic paper. We use LongBench [13] dataset which contains the summarization task [^4] We capped the input lengths in LongBench because OPT’s absolute positional embedding only supports a maximum length of 2048.. As shown in Figure 7, LongBench has much longer input lengths than the other two datasets. So we set a loose TTFT SLO but require a stringent TPOT.
@@
Metrics. We use SLO attainment as the major evaluation metric. Under a specific SLO attainment goal (say, 90%), we are concerned with two things: the maximum per-GPU goodput and the minimal SLO the system can handle. We are particularly interested in an SLO attainment of 90% (indicated by the vertical lines in all curve plots), but will also vary the rate and latency requirements to observe how the SLO attainment changes. We also include the results in the Appendix for an SLO attainment of 99% to show the system performance under a more stringent SLO attainment target.
Baselines. We compare DistServe to two baseline systems:
@@
  • vLLM [32]: vLLM is a representative LLM serving system widely used in both academia and industry. It supports continuous batching [54] to increase throughput and paged-attention [32] to reduce memory fragmentation during KV cache allocation. However, it colocates the prefill and decoding computation to maximize the overall system throughput and struggles to meet the latency requirements cost-efficiently. Since vLLM only supports intra-op parallelism, we follow previous work [32] to set intra-op equals 1, 4, and 8 for the three OPT models, respectively.
  • DeepSpeed-MII [3]: DeepSpeed Model Implementations for Inference (MII) supports chunked-prefill by decomposing long prompts into smaller chunks and composing with short prompts to exactly fill a target token budget. It mitigates but cannot eliminate the prefill-decoding interference caused by the long prefill job. We set its intra-op the same as vLLM for OPT-13B and OPT-66B for a fair comparison. However, DeepSpeed-MII cannot serve OPT-175B whose $vocab\_size=50272$ because its underlying kernel implementation requires $vocab\_size/intra\_op$ is a multiple of 8 where intra-op equals 8 does not satisfy. Setting intra-op equals 4 can satisfy this requirement but will cause the out-of-memory issue.
@@ 6.2 End-to-end Experiments
In this Section, we compare the end-to-end performance of DistServe against the baselines on real application datasets.
@@
Chatbot. We evaluate the performance of DistServe on the chatbot application for all three OPT models. The first row of Figure 8 illustrates that when we gradually increase the rate, more requests will violate the latency requirements and the SLO attainment decreases. The vertical line shows the maximum per-GPU rate the system can handle to meet latency requirements for over 90% of the requests.
@@
On the ShareGPT dataset, DistServe can sustain 2.0×2.0$\times$2.0 ×4.6×4.6$\times$4.6 × higher request rate compared to vLLM. This is because DistLLM eliminates the prefill-decoding interference through disaggregation. Two phases can optimize their own objectives by allocating different resources and employing tailored parallelism strategies. Specifically, by analyzing the chosen placement strategy [^5] All the placements chosen by DistServe can be found in Appendix B. for 175B, we find the prefill instance has inter-op = 3, intra-op = 3; and the decoding instance has inter-op = 3, intra-op = 4. Under this placement, DistServe can effectively balance the load between the two instances on ShareGPT, meeting latency requirements at the lowest cost. This non-trivial placement strategy is challenging to manually find, proving the effectiveness of the algorithm. In the case of vLLM, collocating prefill and decoding greatly slows down the decoding phase, thereby significantly increasing TPOT. Due to the stringent TPOT requirements of chatbot applications, although vLLM meets the TTFT SLO for most requests, the overall SLO attainment is dragged down by a large number of requests that violate the TPOT SLO. Compared to DeepSpeed-MII, DistServe can sustain 1.6×1.6$\times$1.6 ×7.4×7.4$\times$7.4 × higher request rate. DeepSpeed-MII shows better performance on larger models because the prefill job is larger and chunked-prefill mitigates the interference to some extent. However, due to the reasons discussed in §2.3, chunked prefill is slower than full prefill, so it struggles to meet the TTFT SLO as a sacrifice for better TPOT.
@@
The second row of Figure 8 indicates the robustness to the changing latency requirements of the two systems. We fix the rate and then linearly scale the two latency requirements in Table 1 simultaneously using a parameter called SLO Scale. As SLO Scale decreases, the latency requirement is more stringent. We aim to observe the most stringent SLO Scale that the system can withstand while still achieving the attainment target. Figure 8 shows that DistServe can achieve 1.8×1.8$\times$1.8 ×3.2×3.2$\times$3.2 × more stringent SLO than vLLM and 1.7×1.7$\times$1.7 ×1.8×1.8$\times$1.8 × more stringent SLO than DeepSpeed-MII, thus providing more engaging service quality to the users.
@@
Code completion. Figure 9(a) shows the performance of DistServe on the code completion task when serving OPT-66B. DistServe can sustain 5.7×5.7$\times$5.7 × higher request rate and 1.4×1.4$\times$1.4 × more stringent SLO than vLLM. Compared to DeepSpeed-MII, DistServe can sustain 1.6×1.6$\times$1.6 × higher request rate and 1.4×1.4$\times$1.4 × more stringent SLO. As a real-time coding assistant, the code completion task demands lower TTFT than chatbot, this leads to both systems ultimately being constrained by the TTFT requirement. However, in comparison, by eliminating the interference of the decoding jobs and automatically increasing intra-operation parallelism in prefill instances through the searching algorithm, DistServe reduces the average latency of the prefill jobs, thereby meeting the TTFT requirements of more requests.
@@
Summarization. Figure 9(b) shows the performance of DistServe on the summarization task when serving OPT-66B. DistServe achieves 4.3×4.3$\times$4.3 × higher request rate and 12.6×12.6$\times$12.6 × more stringent SLO than vLLM. Compared to DeepSpeed-MII, DistServe achieves 1.8×1.8$\times$1.8 × higher request rate and 2.6×2.6$\times$2.6 × more stringent SLO. The requests sampled from LongBench dataset have long input lengths, which brings significant pressure to the prefill computation. However, due to the loose requirement of TTFT for the summarization task, the TPOT service quality becomes particularly important. Since vLLM collocates prefill and decoding phases, it experiences a greater slowdown in the decoding phase with long prefill jobs and fails to meet the TPOT requirement.
The results above are all under the 90% SLO attainment target. We observe that DistServe can have better performance under a more stringent attainment target (say, 99%) and present the results in Appendix C.
@@
Figure 10: Left: Latency breakdown when serving OPT-175B on ShareGPT dataset with DistServe. Right: The CDF function of KV Cache transmission time for three OPT models.
@@ 6.3 Latency Breakdown
Figure 10(a) shows the latency breakdown for the OPT-175B models on the ShareGPT dataset. We chose OPT-175B because the KV Cache transmission is more demanding for larger models. In fact, even for OPT-175B, the KV Cache transmission only accounts for less than 0.1% of the total latency. Even by examining the CDF of the absolute transmission time shown in Figure 10(b), we observe that over 95% of requests experience a delay of less than 30ms, despite our testbed having only limited cross-node bandwidth. This is due to the algorithm described in §4.2, where we require the prefill and decoding instance to maintain the same stage on one machine, enabling the use of intra-node NVLINK bandwidth for transmission, thus significantly reducing transmission delay.
@@ 6.4 Ablation Studies
Results. Figure 11 shows the performance of the four systems when serving OPT-66B on the ShareGPT dataset. "vLLM++" has the same performance as "vLLM" because we find the default parallelism setting (intra-op=4) has the best per-GPU goodput. This further demonstrates the importance of disaggregation. The interference between the prefill and decoding phases significantly reduces the potential performance improvement through adjusting parallelism. In contrast, "DistLLM-High" can achieve further improvements over "DistLLM-Low" because it is not constrained by the deployment constraint that the prefill and decoding instance on one node should share the same model stage. Through disaggregation, we can use tailored parallelism strategies for prefill and decoding instances and optimize their targets without the coupling effects.